# DVMNet: Computing Relative Pose for Unseen Objects Beyond Hypotheses

### v1

@@ 3.2 Motivation
Drawing inspiration from the success of pixel-pixel correspondences in image matching [45, 49, 26] a natural approach to avoiding the use of pose hypotheses would be to compute the relative object pose based on 2D correspondences. However, recent studies [18, 52] have observed that such an image-matching strategy is unreliable in the scenario of object pose estimation. We trace this limitation back to the fact that image-matching methods are not fully differentiable w.r.t. the pose. Specifically, some approaches [45, 48, 46] encode a notion of consistency among the pixel-pixel correspondences utilizing the essential matrix. However, computing the pose from the essential matrix leads to multiple solutions [13]. The pose estimation is thus detached from the learning process as a post-processing step. Notably, in the context of object pose estimation, the pre-generated correspondences tend to be unreliable in the presence of challenges such as large object pose differences and textureless objects. Therefore, the isolated pose estimation step in the two-stage design becomes less effective.
@@
To address this issue, we propose to lift the input images to voxelized 3D representations [40] and perform the matching process in 3D space. Therefore, the computation of the relative object pose from the resulting voxel-voxel matches becomes a differentiable operation. This characteristic enables us to directly supervise the entire framework with the actual quantity we aim to predict, i.e., the relative object pose. Subsequently, we elaborate on the steps involved in the presented hypothesis-free mechanism.
3.3 Image Voxelization
@@
To achieve the pose estimation from voxel-voxel correspondences with only RGB images as input, we first need to lift each RGB image to a set of 3D voxels. To enable such a voxelization, we introduce an autoencoder network depicted in Fig. 2, which includes a 2D-3D encoder and a 3D-2D decoder. Specifically, we employ a pretrained vision transformer [41] to convert the query and reference images to 2D feature embeddings denoted as $\mathbf{F}_{q}$ and $\mathbf{F}_{r}$, respectively. Considering the difficulty of lifting 2D images to 3D representations, we incorporate a cross-attention module to capture cross-view 3D information. We take the feature embedding $\mathbf{F}_{q}$ as an example (a symmetric process is carried out for $\mathbf{F}_{r}$). The cross-attention module [41] is defined as
@@
Benefiting from such a 3D-aware encoding process, we voxelize the image feature embeddings via a simple reshaping process. Note that, to facilitate the robust pose estimation that will be introduced in Sec. 3.5, we predict an objectness score for each voxel, which reflects the significance of the voxel to the relative object pose estimation. Therefore, the actual reshaping process is conducted as $\mathbb{R}^{C\times H\times W}\rightarrow\mathbb{R}^{(C^{{}^{\prime}}+1)\times D\times H\times W}$, where $C=(C^{{}^{\prime}}+1)\times D$. As shown in Fig. 2, we denote the resulting 3D objectness maps and 3D volumes as $\mathbf{O}_{q},\mathbf{O}_{r}\in\mathbb{R}^{1\times D\times H\times W}$ and $\mathbf{V}_{q},\mathbf{V}_{r}\in\mathbb{R}^{C^{{}^{\prime}}\times D\times H\times W}$, respectively.
As our approach does not rely on object segmentation, the learned voxel representations might be affected by the background of the query and reference images. To alleviate this issue, we introduce an object-aware decoding process over $\mathbf{V}_{q}$ and $\mathbf{V}_{r}$. Concretely, $\mathbf{V}_{q}$ and $\mathbf{V}_{r}$ are projected to 2D space by aggregating the voxels along the depth direction as $\mathbb{R}^{C^{{}^{\prime}}\times D\times H\times W}\rightarrow\mathbb{R}^{C^{*}\times H\times W}$, where $C^{*}=C^{{}^{\prime}}\times D$. The resulting 2D feature embeddings are then fed into a decoder that contains several self-attention modules [10] from which the object images $\hat{\mathbf{I}}_{q}$ and $\hat{\mathbf{I}}_{r}$ without background are produced. The object masks $\hat{\mathbf{M}}_{q}$ and $\hat{\mathbf{M}}_{r}$ are additionally predicted to provide auxiliary information that benefits the following robust object pose estimation.
@@
3.4 Object Pose from Deep Voxel Matching
Figure 3: Computing relative object pose from 3D voxels. The feature similarities of $\mathbf{V}_{q}$ and $\mathbf{V}_{r}$ are computed by our deep voxel matching module, resulting in a score matrix $\mathbf{S}$. A soft assignment is performed based on $\mathbf{S}$ over the query object mask $\hat{\mathbf{M}}_{q}$, the 3D objectness map $\mathbf{O}_{q}$, and the 3D coordinates $\mathbf{X}_{q}$. The aligned query and reference voxels are then fed into a Weighted Closest Voxel (WCV) algorithm that estimates the relative object pose in a robust and end-to-end manner.
According to multi-view geometry [4, 13, 38], the relative object pose can be computed by solving a least-squares problem expressed in terms of voxel-voxel correspondences. Specifically, as we focus on 3D rotation estimation, the least-squares problem is formulated as
@@ 3.4 Object Pose from Deep Voxel Matching
where $\mathbf{x}_{r}^{i}\in\mathbf{X}_{r}$ and $\mathbf{x}_{q}^{i}\in\mathbf{X}_{r}$ stand for the 3D coordinates of the $i$-th reference and query voxels, respectively. The optimal $\Delta{\hat{\mathbf{R}}}$ is then determined as
@@
Inspired by the studies [38, 16, 19] showing that object pose estimation benefits from the end-to-end training, we carry out the alignment in a differentiable fashion. As illustrated in Fig. 3, the alignment is conducted based on a deep voxel matching module. Specifically, we compute a score matrix $\mathbf{S}$ whose entry $s_{ij}$ indicates the cosine similarities between two voxels as
@@
3.5 Weighted Closest Voxel Algorithm
Figure 4: Illustration of the voxel weights. Each colored dot indicates the voxel position in 3D space. The green dots in the middle and right cubes represent the voxels with larger weights. The voxel weights in the middle cube are computed based on the replicated object mask, while the weights in the right cube are updated by integrating the 3D objectness map with the object mask.
Note that our task differs from standard point cloud registration [28, 38, 1], which typically operates on 3D point clouds sampled from 3D object meshes [6] or captured using specific sensors [8]. Here, by contrast, we work with 3D volumes lifted from 2D images, and some voxels could thus be outliers since the corresponding 2D image patches may depict nuisances such as the background. The presence of these outliers may impact the accuracy of the relative object pose estimated from the voxel matches. To address this challenge, we introduce a weighted closest voxel algorithm that enables robust relative object pose estimation.
Concretely, the objective is to mitigate the effect of unreliable voxel matches. We thus incorporate a weight vector into the pose estimation process, modifying Eq. 10 as
@@ 3.5 Weighted Closest Voxel Algorithm
where $\omega^{i}\in(0,1)$ denotes the weight of the $i$-th voxel pair. This makes the subsequent relative object pose estimation aware of the reliability of each voxel pair. We determine the weight vector by utilizing both the object mask and voxel objectness information produced by the encoder network. Specifically, we first replicate $\mathbf{\hat{M}}_{q}$ and $\mathbf{\hat{M}}_{r}$ $D$ times along the depth dimension, which creates pseudo 3D masks, $\mathbf{\hat{M}}_{q}^{*},\mathbf{\hat{M}}_{r}^{*}\in\mathbb{R}^{1\times D\times H\times W}$. These pseudo 3D masks contribute to alleviating the influence of voxels that depict the background. The weight of each voxel pair is then determined as
@@
Fig. 4 provides an example of the estimated voxel weights. The dots denote the 3D voxel positions and the voxels assigned with larger weights are colored in green within the middle and right cubes. In the right cube, the green dots roughly depict a 3D surface that corresponds to the object visible in the 2D image. Note that our DVMNet is trained without relying on ground-truth 3D object models. This observation thus demonstrates that the voxels crucial in determining the relative object pose are aware of the 3D object shape information. The complete DVMet is trained end-to-end with a loss function defined as $L=L_{ae}+L_{pose}$ with
@@
where $\Delta\mathbf{R}^{gt}$ is the ground-truth relative object pose, and $q(\cdot)$ is a function that converts a rotation matrix to a 6D continuous representation [53].
@@ 4 Experiments
  Method Ball Book Couch Frisbee Hotdog Kite Remote Sandwich Skateboard Suitcase Mean
SuperGlue [26] 83.55 71.02 45.14 68.67 88.74 56.46 78.58 73.64 72.14 76.74 71.47
@@
DVMNet 28.31 21.98 19.01 23.23 21.45 17.50 11.39 19.63 20.14 16.85 19.95
Table 1: Experimental results on CO3D [25]. We report the angular errors of the estimated relative object rotations. All testing object categories were unseen during training. The best results are shown in bold fonts.
@@ 4.1 Implementation Details
In the presented autoencoder, we use 3 cross-attention modules in the 2D-3D encoder and 3 self-attention modules in the 3D-2D decoder. In the relative object pose estimation module, we normalize the 3D coordinates of the voxels to an interval of $[-1,1]$ with a mean of $\mathbf{0}$. We set the temperatures $\tau$ and $\lambda$ in Eq. 13 and Eq. 15 to $0.1$ and $1.0$, respectively. We train our network in an end-to-end fashion on an A100 GPU, employing the AdamW [22] optimizer with a batch size of 64 and a learning rate of $10^{-5}$. Since we focus on predicting the relative object rotation, we crop the object from the RGB image by utilizing the ground-truth object bounding box, following the implementation in [42, 51, 47, 18, 52].
4.2 Relative Object Pose Estimation on CO3D
@@
We first evaluate our method on the CO3D dataset [25], which has been commonly utilized in the literature [47, 18, 52]. This dataset contains 18,619 video sequences that depict 51 object categories. To evaluate the generalization ability of the network to unseen objects, we follow the setting in [47], training the network on 41 object categories and testing it on the other 10 categories. The performance is measured by the mean angular error $e\in[0^{\circ},180^{\circ}]$ of the estimated relative object rotation, which is defined as
$e=\text{arccos}\left(\frac{\text{tr}(\Delta{\hat{\mathbf{R}}^{\text{T}}}\Delta{\mathbf{R}^{gt}})-1}{2}\right).$ (17)
@@
We compare our approach with state-of-the-art techniques including image-matching methods, SuperGlue [26], LoFTR [31], and ZSP [12], hypothesis-based methods, RelPose [47], RelPose++ [18], and 3DAHV [52], and a direct regression method implemented in [18].
@@
As reported in Table 1, DVMNet delivers superior relative pose estimation performance for unseen objects, outperforming both the image-matching and hypothesis-based competitors by at least $8.49^{\circ}$ in terms of mean angular error. To shed more light on the robustness of the evaluated approaches, we categorize the testing image pairs into different groups according to the corresponding angular errors observed when applying a particular relative object pose estimation method. We count the number of image pairs in each group and show the results in Fig. 5. Our method results in a higher number of image pairs with smaller angular errors. More importantly, as highlighted by the red dash box in Fig. 5, both image-matching and hypothesis-based methods exhibit large angular errors for some image pairs. By contrast, our DVMNet results in fewer failure instances, thus demonstrating better robustness.
@@
Figure 5: Histogram depicting the distribution of angular errors. The image pairs in the testing set are divided into distinct groups based on the angular errors obtained by a specific pose estimation approach. Each bar in the histogram represents the count of image pairs within a particular group. Our DVMNet yields much fewer unnaturally large errors than image-matching and hypothesis-based methods.
@@
Furthermore, as argued in Sec. 3, our hypothesis-free strategy is more efficient than the hypothesis-based techniques in relative object pose estimation. We thus assess their computational cost, utilizing the multiply-accumulate operations (MACs). For hypothesis-based methods, all sampled hypotheses are processed in parallel. The results shown in Fig. 1 indicate the benefits of our hypothesis-free DVMNet, which requires considerably fewer MACs than the hypothesis-based competitors. To further substantiate this advantage, we provide detailed results in Fig. 6, where the hypothesis-based methods are evaluated with the number of pose samples varying from 1,000 to 500,000. Note that for 3DAHV, the maximum number is 100,000 because of our computational resource constraints. As shown in Fig. 6, one can enhance the efficiency of the hypothesis-based methods by reducing the number of samples. However, this efficiency gain comes at the cost of sacrificing pose estimation accuracy. By contrast, our method achieves a good trade-off between efficiency and pose estimation accuracy. We also evaluate the time consumption on an A100 GPU. On average, DVMNet processes a pair of images in 23ms. Despite benefiting from parallel estimation, RelPose++ and 3DAHV still cost 29ms and 35ms, respectively, which are slower than our method.
@@
Figure 6: Comparison with hypothesis-based methods. We measure the computational cost as multiply-accumulate operations (MACs). The results for hypothesis-based methods are shown with varying numbers of pose samples, ranging from 1,000 to 500,000. For 3DAHV, we set the maximum number to be 100,000 due to the computational resource constraints.
LINEMOD
SuperGlue [26]
LoFTR [31]
ZSP [12] Regress [18] RelPose [47] RelPose++ [18] 3DAHV [52] DVMNet
Cat
67.28
88.06
79.61 54.21 53.72 47.77 50.99 31.70
Ben.
58.52
70.80
74.07 52.03 62.32 44.67 38.16 34.00
Cam.
58.11
87.13
79.65 51.04 59.91 44.31 41.92 33.18
Dri.
65.16
78.85
76.35 52.83 57.61 47.95 32.65 46.29
Duck
74.90
97.63
83.43 55.44 55.15 48.65 44.03 38.91
Mean
64.79
84.49
78.62 53.11 57.75 46.67 41.55 36.82
Objaverse
SuperGlue [26]
LoFTR [31]
ZSP [12] Regress [18] RelPose [47] RelPose++ [18] 3DAHV [52] DVMNet
Mean
102.40
134.05
107.20 55.90 80.39 33.49 28.11 20.19
@@
Table 2: Experimental results on the GROP benchmark [52]. The methods are evaluated in terms of angular error on the LINEMOD [15] and Objaverse [9] datasets. The testing data comprises 5 objects from LINEMOD and 128 objects from Objaverse. All images containing these objects are omitted from the training set.
4.3 Relative Object Pose Estimation on GROP

### v2

@@ 3.3.1 Motivation
Drawing inspiration from the success of pixel-pixel correspondences in image matching [46, 16, 47], a natural approach to avoiding the use of rotation hypotheses would be to compute the relative object rotation based on 2D correspondences. However, recent studies [13, 14] have observed that such an image-matching strategy is unreliable in the scenario of object pose estimation. We trace this limitation back to the fact that image-matching methods are not fully differentiable w.r.t. the rotation. Specifically, some approaches [46, 48, 49] encode a notion of consistency among the pixel-pixel correspondences utilizing the essential matrix. However, computing the rotation from the essential matrix leads to multiple solutions [15]. Rotation estimation is thus detached from the learning process as a post-processing step. Notably, in the context of object rotation estimation, those pre-generated correspondences tend to be unreliable in the presence of challenges such as large object rotation differences and textureless objects. Therefore, the isolated rotation estimation step in the two-stage design becomes less effective.
@@
To address this issue, we propose to lift the input images to voxelized 3D representations [50] and perform the matching process in 3D latent space. Therefore, the computation of the relative object rotation from the resulting voxel-voxel matches becomes a differentiable operation. This characteristic enables us to directly supervise the rotation estimation module with the actual quantity we aim to predict, i.e., the relative object pose. Below, we elaborate on the steps involved in the presented hypothesis-free mechanism.
3.3.2 Image Voxelization
@@
To achieve object rotation estimation from voxel-voxel correspondences with only RGB images as input, we first need to lift each RGB image to a set of 3D voxels. To enable such a voxelization, we introduce an autoencoder network depicted in Fig. 4, which includes a 2D-3D encoder and a 3D-2D decoder. Specifically, we employ a pretrained vision transformer [51] to convert the query and reference images to 2D feature embeddings denoted as $\mathbf{F}_{q}$ and $\mathbf{F}_{r}$, respectively. Considering the difficulty of lifting 2D images to 3D representations, we incorporate a cross-attention module to capture cross-view 3D information. We take the feature embedding $\mathbf{F}_{q}$ as an example (a symmetric process is carried out for $\mathbf{F}_{r}$). The cross-attention module [51] is defined as
@@
Benefiting from such a 3D-aware encoding process, we voxelize the image feature embeddings via a simple reshaping process. Note that, to facilitate the robust rotation estimation that will be introduced in Section 3.3.4, we predict an objectness score for each voxel, which reflects the significance of the voxel to the relative object rotation estimation. Therefore, the actual reshaping process is conducted as $\mathbb{R}^{C\times H\times W}\rightarrow\mathbb{R}^{(C^{{}^{\prime}}+1)\times D\times H\times W}$, where $C=(C^{{}^{\prime}}+1)\times D$. As shown in Fig. 4, we denote the resulting 3D objectness maps and 3D volumes as $\mathbf{O}_{q},\mathbf{O}_{r}\in\mathbb{R}^{1\times D\times H\times W}$ and $\mathbf{V}_{q},\mathbf{V}_{r}\in\mathbb{R}^{C^{{}^{\prime}}\times D\times H\times W}$, respectively. Since our approach does not rely on object segmentation, the learned voxel representations may be affected by the background of the query and reference images. To alleviate this issue, we introduce an object-aware decoding process over $\mathbf{V}_{q}$ and $\mathbf{V}_{r}$. Concretely, $\mathbf{V}_{q}$ and $\mathbf{V}_{r}$ are projected to 2D space by aggregating the voxels along the depth direction as $\mathbb{R}^{C^{{}^{\prime}}\times D\times H\times W}\rightarrow\mathbb{R}^{C^{*}\times H\times W}$, where $C^{*}=C^{{}^{\prime}}\times D$. The resulting 2D feature embeddings are then fed into a decoder that contains several self-attention modules [53] from which the object images $\hat{\mathbf{I}}_{q}$ and $\hat{\mathbf{I}}_{r}$ without background are produced. The object masks $\hat{\mathbf{M}}_{q}$ and $\hat{\mathbf{M}}_{r}$ are additionally predicted to provide auxiliary information that benefits the following robust object rotation estimation.
@@
3.3.3 Object Rotation from Deep Voxel Matching
Figure 5: Computing relative object rotation from 3D voxels. The feature similarities of $\mathbf{V}_{q}$ and $\mathbf{V}_{r}$ are computed, which results in a score matrix $\mathbf{S}$. A soft assignment is performed based on $\mathbf{S}$ over the query object mask $\hat{\mathbf{M}}_{q}$, the 3D objectness map $\mathbf{O}_{q}$, and the 3D coordinates $\mathbf{X}_{q}$. The aligned query and reference voxels are then fed into a Weighted Closest Voxel (WCV) algorithm that estimates the relative object rotation in a robust and end-to-end manner.
According to multi-view geometry [54, 15, 55], relative object rotation can be computed by solving a least-squares problem expressed in terms of voxel-voxel correspondences. Specifically, the least-squares problem is formulated as
@@ 3.3.3 Object Rotation from Deep Voxel Matching
where $\mathbf{x}_{r}^{i}\in\mathbf{X}_{r}$ and $\mathbf{x}_{q}^{i}\in\mathbf{X}_{r}$ stand for the 3D coordinates of the $i$-th reference and query voxels, respectively. The coordinates are normalized to be zero-centered and unit-scale. The optimal $\Delta{\hat{\mathbf{R}}}$ is then determined as
@@
Inspired by the studies [55, 56, 34] showing that object pose estimation benefits from end-to-end training, we carry out the alignment in a differentiable fashion. As illustrated in Fig. 5, the alignment is conducted based on a deep voxel matching module. Specifically, we compute a score matrix $\mathbf{S}$ whose entry $s_{ij}$ indicates the cosine similarity between two voxels as
@@
3.3.4 Weighted Closest Voxel Algorithm
Figure 6: Illustration of the voxel weights. Each colored dot indicates the voxel position in 3D space. The green dots in the middle and right cubes represent the voxels with larger weights. The voxel weights in the middle cube are computed based on the replicated object mask. The weights in the right cube are updated by integrating the 3D objectness map with the object mask.
Note that our task differs from standard point cloud registration [57, 55, 58], which typically operates on 3D point clouds sampled from 3D object meshes [59] or captured using specific sensors [60]. Here, by contrast, we work with 3D volumes lifted from 2D images, and some voxels could thus be outliers since the corresponding 2D image patches may depict nuisances such as the background. The presence of these outliers may impact the accuracy of the relative object rotation estimated from the voxel matches. To address this challenge, we introduce a weighted closest voxel algorithm that enables robust relative object rotation estimation.
Concretely, the objective is to mitigate the effect of unreliable voxel matches. We thus incorporate a weight vector into the rotation estimation process, modifying Eq. 13 as
@@ 3.3.4 Weighted Closest Voxel Algorithm
where $\omega^{i}\in(0,1)$ denotes the weight of the $i$-th voxel pair. This makes the subsequent relative object rotation estimation aware of the reliability of each voxel pair. We determine the weight vector by utilizing both the object mask and voxel objectness information produced by the encoder network. Specifically, we first replicate $\mathbf{\hat{M}}_{q}$ and $\mathbf{\hat{M}}_{r}$ $D$ times along the depth dimension, which creates pseudo 3D masks, $\mathbf{\hat{M}}_{q}^{*},\mathbf{\hat{M}}_{r}^{*}\in\mathbb{R}^{1\times D\times H\times W}$. These pseudo 3D masks contribute to alleviating the influence of voxels that depict the background. The weight of each voxel pair is then determined as
@@
Fig. 6 provides an example of the estimated voxel weights. The dots denote the 3D voxel positions and the voxels assigned with larger weights are colored in green within the middle and right cubes. In the right cube, the green dots roughly depict a 3D surface that corresponds to the object visible in the 2D image. Note that our rotation estimation network is trained without relying on ground-truth 3D object models. This observation thus demonstrates that the voxels that are crucial in determining the relative object rotation are aware of the 3D object shape information. The complete rotation estimation module is trained end-to-end with a loss function defined as $L=L_{ae}+L_{pose}$ with
@@
where $\Delta\mathbf{R}^{gt}$ is the ground-truth relative object rotation, and $q(\cdot)$ is a function that converts a rotation matrix to a 6D continuous representation [37].
@@ 4 Experiments
TABLE I: Relative object rotation estimation on CO3D [23]. We report the angular errors of the estimated relative object rotations. All testing object categories were unseen during training. The best results are shown in bold fonts.
  Method Ball Book Couch Frisb. Hotd. Kite Remot. Sandw. Skate. Suitc. Mean
SG [16] 83.55 71.02 45.14 68.67 88.74 56.46 78.58 73.64 72.14 76.74 71.47
@@
DVMNet++ 28.31 21.98 19.01 23.23 21.45 17.50 11.39 19.63 20.14 16.85 19.95
@@ 4.1 Implementation Details
In the presented autoencoder, we use 3 cross-attention modules in the 2D-3D encoder and 3 self-attention modules in the 3D-2D decoder. In the relative object rotation estimation module, we normalize the 3D coordinates of the voxels to an interval of $[-1,1]$ with a mean of $\mathbf{0}$. We set the temperatures $\tau$ and $\lambda$ in Eq. 16 and Eq. 18 to $0.1$ and $1.0$, respectively. We train our network on an A100 GPU, employing the AdamW [61] optimizer with a batch size of 64 and a learning rate of $10^{-5}$. We crop the object from the query image using the ground-truth object bounding box during the training stage, following the implementation in [12, 13, 14]. We replace the ground truth with the bounding box predicted by the proposed open-set object detector at test time.
4.2 Relative Object Rotation Estimation on CO3D
@@
We first evaluate our method on the CO3D dataset [23], which has been commonly utilized in the literature [12, 13, 14]. This dataset contains 18,619 video sequences that depict 51 object categories. To evaluate the generalization ability of the network to unseen objects, we follow the setting in [12], training the network on 41 object categories and testing it on the other 10 categories. The performance is measured by the mean angular error $err_{R}\in[0^{\circ},180^{\circ}]$ of the estimated relative object rotation, which is defined as
$err_{R}=\text{arccos}\left(\frac{\text{tr}(\Delta{\hat{\mathbf{R}}^{T}}\Delta{\mathbf{R}_{gt}})-1}{2}\right).$ (20)
@@
We compare our approach with state-of-the-art techniques including image-matching methods, SuperGlue (SG) [16], LoFTR [17], and ZSP [36], hypothesis-based methods, RelPose [12], RelPose++ [13], and 3DAHV [14], a diffusion-based method, PoseDiffusion [38], and a direct regression method implemented in [13]. Note that since we focus on the evaluation of relative rotation estimation, we use the ground-truth object bounding box to locate the object in the query image and only utilize a single reference image. We maintain this setting across all evaluated methods to ensure a fair comparison.
@@
As reported in Table I, DVMNet++ delivers superior relative rotation estimation performance for unseen objects, outperforming both the image-matching and hypothesis-based competitors by at least $8.49^{\circ}$ in terms of mean angular error. To shed more light on the robustness of the evaluated approaches, we categorize the testing image pairs into different groups according to the corresponding angular errors observed when applying a particular relative object rotation estimation method. We count the number of image pairs in each group and show the results in Fig. 7. Our method results in a higher number of image pairs with smaller angular errors. More importantly, as highlighted by the red dashed box in Fig. 7, both image-matching and hypothesis-based methods exhibit large angular errors for some image pairs. By contrast, our DVMNet++ yields fewer failure instances, thus demonstrating better robustness.
@@
Figure 7: Histogram depicting the distribution of rotation errors. The image pairs in the testing set are divided into distinct groups based on the angular errors obtained by a specific rotation estimation approach. Each bar in the histogram represents the count of image pairs within a particular group. Our DVMNet++ yields much fewer unnaturally large errors than image-matching and hypothesis-based methods.
TABLE II: Time consumption. We evaluate the speed on an A100 GPU. The average time consumption per image pair is reported. For hypothesis-based approaches, the rotation hypotheses are processed in parallel.
  RelPose++ PoseDiffusion 3DAHV DVMNet++
29ms 5584ms 35ms 23ms
@@
Furthermore, as argued in Section 3, our hypothesis-free strategy is more efficient than the hypothesis-based techniques in relative object rotation estimation. We thus assess their computational cost, utilizing the multiply-accumulate operations (MACs). For hypothesis-based methods, all sampled hypotheses are processed in parallel. The results shown in Fig. 1 indicate the benefits of our hypothesis-free DVMNet++, which requires considerably fewer MACs than the hypothesis-based competitors. To further substantiate this advantage, we provide detailed results in Fig. 8, where the hypothesis-based methods are evaluated with the number of rotation samples varying from 1,000 to 500,000. Note that for 3DAHV, the maximum number is 100,000 because of our computational resource constraints. As shown in Fig. 8, one can enhance the efficiency of the hypothesis-based methods by reducing the number of samples. However, this efficiency gain comes at the cost of sacrificing rotation estimation accuracy. By contrast, our method achieves a good trade-off between efficiency and rotation estimation accuracy. We also evaluate the time consumption on an A100 GPU and the results are reported in Table II. On average, DVMNet++ processes a pair of images in 23ms. Despite benefiting from parallel estimation, the hypothesis-based methods RelPose++ and 3DAHV are still slower than our method.
@@
Figure 8: Comparison with hypothesis-based methods. We measure the computational cost as multiply-accumulate operations (MACs). The results for hypothesis-based methods are shown with varying numbers of rotation samples, ranging from 1,000 to 500,000. For 3DAHV, we set the maximum number to be 100,000 due to the computational resource constraints.
@@
TABLE III: Relative object rotation estimation on the GROP benchmark [14]. The methods are evaluated in terms of angular error on the LINEMOD [18] and Objaverse [24] datasets. The testing data comprises 5 objects from LINEMOD and 128 objects from Objaverse. All images containing these objects are omitted from the training set.
  LINEMOD SG [16] LoFTR [17] ZSP [36] Regress [13] RelPose [12] RelPose++ [13] 3DAHV [14] DVMNet++
Cat 67.28 88.06 79.61 54.21 53.72 47.77 50.99 31.70
Ben. 58.52 70.80 74.07 52.03 62.32 44.67 38.16 34.00
Cam. 58.11 87.13 79.65 51.04 59.91 44.31 41.92 33.18
Dri. 65.16 78.85 76.35 52.83 57.61 47.95 32.65 46.29
Duck 74.90 97.63 83.43 55.44 55.15 48.65 44.03 38.91
Mean 64.79 84.49 78.62 53.11 57.75 46.67 41.55 36.82
  Objaverse SG [16] LoFTR [17] ZSP [36] Regress [13] RelPose [12] RelPose++ [13] 3DAHV [14] DVMNet++
Mean 102.40 134.05 107.20 55.90 80.39 33.49 28.11 20.19
4.3 Relative Object Rotation Estimation on GROP